A boat travels 24 mi upstream in 3 hrs. then it takes 3 hrs to go downstream at 36 mi downstream, Find the speed of boat in still water and the speed of the current. (using two unknowns, write two equations and solve.)
>>>>>>>>>>>>>THANKS IN ADVANCE

Answer by stanbon(57984)

A boat travels 24 mi upstream in 3 hrs. then it takes 3 hrs to go downstream at 36 mi downstream, Find the speed of boat in still water and the speed of the current. (using two unknowns, write two equations and solve.)
----------------
Let still water speed be "b"; Let current speed be "c"
-------------------
Upstream DATA:
Distance = 24 miles ; Time = 3 hrs ; Rate = b-c=d/t = 8 mph
--------------
Downstream DATA:
Distance = 36 miles ; Time = 3 hrs ; Rate = b+c=d/t = 12 mph
-------------------
EQUATIONS:
b+c=12
b-c=8
--------
Add to get:
2b = 20
b = 10 mph (boat speed in still water)
----
Substitute to solve for "c":
10+c = 12
c = 2 mph (current speed)
==============
Cheers,
Stan H.